OpenStudy (anonymous):

SOMEONE PLEASE HELP I WILL GIVE A MEDAL Solve for x. y = mx + b and what is the varible

OpenStudy (muzzack):

mx=b+y x=b/m+y/m

OpenStudy (muzzack):

answer choices?

OpenStudy (anonymous):

i just need to know how to work it out